### 01 Feb 2024 v1

@@ III Excess at 95 GeV
The value for $\mu_{LEP}^{bb}$ can be found in 98LEP1 ; 98LHC with the method introduced in 98LEP2 . $h_{1}$ is the Higgs which mass range around $\sim$ 96 GeV and $h_{2}$ is the 125 GeV Higgs boson in the following. In the feamework of the $\mu\nu$SSM we use $\mu_{NP}^{bb}$ to describe the signal strength, the expression for $\mu_{NP}^{bb}$ can be approximated as 98LEP3
@@
One can found the SM branching ratios $Br^{SM}$ in Ref. PDG1 , $\Gamma$ is the decay widths. The couplings are normalized to the SM-prediction of a Higgs boson of the same mass. $C_{h_{1}}$ is the coupling of $h_{1}$ and gauge boson, $C_{h_{1}u\bar{u}}$ and $C_{h_{1}d\bar{d}}$ are the couplings of $h_{1}$ and up- and down-type quarks. The normalized couplings are given in:
$C_{h_{i}d\bar{d}}={{Z_{H}^{i1}}\over{\cos\beta}},\qquad C_{h_{i}u\bar{u}}={{Z_{H}^{i2}}\over{\sin\beta}},\qquad C_{h_{i}VV}=Z_{H}^{i1}\cos\beta+Z_{H}^{i2}\sin\beta,$ (30)
@@
$\mu_{\gamma\gamma}^{CMS}={{\sigma(gg\rightarrow h_{1}\rightarrow\gamma\gamma)}\over{\sigma^{SM}(gg\rightarrow h_{1}\rightarrow\gamma\gamma)}}=0.6\pm 0.2,$ (31)
Recently, ATLAS reported their new results at 95.4 GeV based on full Run 2 data set 95ATLAS , the “di-photon excess” with a signal strength of
@@
Meanwhile, the corresponding CMS result for the “di-photon excess” is given by CMS4
@@
In this work, the approximation of the diphoton rate of the $h_{1}$ can written as 98LEP1 ; 98LEP3 :
@@
$\approx{{|C_{h_{1}u\bar{u}}|^{2}\times|C_{h_{1}\gamma\gamma}|^{2}}\over{|C_{h_{1}d\bar{d}}|^{2}(Br_{h_{1}\rightarrow b\bar{b}}^{\text{SM}}+Br_{h_{1}\rightarrow\tau\bar{\tau}}^{\text{SM}})+|C_{h_{1}u\bar{u}}|^{2}(Br_{h_{1}\rightarrow gg}^{\text{SM}}+Br_{h_{1}\rightarrow c\bar{c}}^{\text{SM}})}}$
@@
$|C_{h_{1}\gamma\gamma}|^{2}={{|{4\over 3}C_{h_{1}t\bar{t}}A_{1/2}(\tau_{t})+C_{h_{1}VV}A_{1}(\tau_{W})|^{2}}\over{|{4\over 3}A_{1/2}(\tau_{t})+A_{1}(\tau_{W})|^{2}}}.$ (36)
with $\tau_{t}={{m_{h_{1}}^{2}}\over{4m_{t}^{2}}}<1$ and $\tau_{t}={{m_{h_{1}}^{2}}\over{4m_{W}^{2}}}<1$, the from factor $A_{1/2}$ and $A_{1}$ are given by Chrr
@@
By using Eq.(29) and Eq.(35), we calculate the two signal strengths.
@@ IV Numerical Results
$\lambda,\tan\beta,\kappa,A_{t},A_{\lambda},A_{\kappa},A_{b},\upsilon_{\nu^{c}}$ (39)
we take $\upsilon_{\nu_{1}^{c}}=\upsilon_{\nu_{2}^{c}}=\upsilon_{\nu_{3}^{c}}$, and we have defined :
@@ IV.1 Mass and Coupling
Figure 1: Values for $m_{h_{2}}$ versus $m_{h_{2}}$ in (a), the normalized couplings $|C_{h_{2}VV}|$ versus $|C_{h_{1}VV}|$ in (b). The input parameters in Tab. 1.
[…]
Figure 2: Values for $m_{h_{1}}$ versus the parameters $\upsilon_{\nu^{c}}$ in (a) and $\kappa$ in (c), values for $m_{h_{2}}$ versus the parameters $\upsilon_{\nu^{c}}$ in (b) and $\kappa$ in (d). We take $\tan\beta=5$ and other parameters in Tab. 1.
In Fig.2, the left column, we can see that $m_{h_{1}}$ is very sensitive to $\upsilon_{\nu^{c}}$. As $\upsilon_{\nu^{c}}$ becomes larger, $m_{h_{1}}$ will rapidly become smaller, on the contrary, $m_{h_{2}}$ will slowly become larger. In order for $m_{h_{1}}$ to be around 95.4 GeV, the value of $\upsilon_{\nu^{c}}$ will not have a wide range. Similar to $\upsilon_{\nu^{c}}$, $m_{h_{1}}$ is also very sensitive to $\kappa$, and with the increase of $\lambda$, a larger $\kappa$ value can be taken to keep $m_{h_{1}}$ near 95.4 GeV. On the contrary, $\kappa$ has little effect on $m_{h_{2}}$, especially when $\lambda$=0.085 or 0.087, $m_{h_{2}}$ will be very stable.
@@
Figure 3: The component of $h_{1}$ (a) and $h_{2}$ (b). The input parameters in Tab.1, and we take $\tan\beta=4.31$.
In Fig. 3, we have showed the component of $h_{1}$ and $h_{2}$, for $h_{1}$, CP-odd $H_{u}$ component and CP-even right-handed sneutrinos component are the main components, and with the increase of $\lambda$, CP-odd $H_{d}$ component will become smaller, while the CP-even right-handed sneutrino component will become larger. While for $h_{2}$, CP-even $H_{u}$ is the main component, as $\lambda$ increases, the CP-even $H_{u}$ component will increase, the CP-even $H_{d}$ component will gradually decrease, and the CP-odd $H_{d}$ component will gradually increase.
Figure 4: The left column is the correlation of $h_{1}$ and the CPV phases, the grey area represents the experimental error of 3 GeV. The right column is the correlation of $h_{2}$ and the CPV phases, the grey area is the theory uncertainty of 1 GeV.
Then, in Fig. 4, we analyze the correlation between the Higgs masses and the CPV phases. For the first row, we take $\lambda=0.09,\tan\beta=6,\kappa=0.315,$ for the second row, $\lambda=0.095,\tan\beta=4.5,\kappa=0.315,$ for the third row, $\lambda=0.098,\tan\beta=4.3,\kappa=0.32$, for the last row, $\lambda=0.085,\tan\beta=3,\kappa=0.04$. Other parameter keep same as Tab.1. We should remark that there are many CPV phase values that can be constrained by the Higgs masses of Fig.7 (a), and here we only select one of them. The signal strengths and the Higgs masses do not increase or decrease all the time, but oscillates like a sine function. In Fig. 4, the first row, when $\phi_{\lambda}$ grown from 4 to 5, the mass range of the “di-photon excess” can fit the experimental constraints, the SM-like Higgs mass can be kept around 125 GeV.
@@ IV.2 Signal Strengths
Figure 5: The left column shows that the signal strength $\mu_{NP}^{bb}$ varies with $\upsilon_{\nu^{c}}$ (a) and $\kappa$ (c), the right column shows that the signal strength $\mu_{NP}^{\gamma\gamma}$ varies with $\upsilon_{\nu^{c}}$ (b) and $\kappa$ (d). We take $\lambda=0.09$ and other parameters in Tab. 1.
@@
The effects of CPV phases is also very obvious. Here, we take $\phi_{A_{t}}$ as an example, we take $\phi_{A_{t}}$ from $-\pi$ to $\pi$. In the first row of Fig. 6, we can see that the minimum value of $h_{1}$ appears to be around 96.2 GeV and varies periodically with $\phi_{A_{t}}$. The maximum value of $h_{1}$ does not exceed 1 GeV larger than the minimum value. Meanwhile, the highest point of $h_{2}$ is at 125.3 GeV, with a maximum value 1.5GeV larger than the minimum value.
The impact of $\phi_{A_{t}}$ on the normalized couplings are relatively small. In the second row of Fig. 6, we can see that the peak of $|C_{h_{1}VV}|$ is close to 0.313, and the highest point of $|C_{h_{2}VV}|$ is slightly less than 0.856. This implies that if the ”95.4 excess” is a new particle, then its coupling with gauge bosons should be much smaller than the coupling of the SM-like Higgs with gauge bosons.
Figure 6: Higgs masses, couplings, and signal strengths are sensitive to CPV phases. Here, we take $\tan\beta$=4.5, $\lambda$=0.086, $\phi_{A_{t}}=[-\pi,\pi]$, other parameters are the same as Tab. 1.
@@
Let us remark that in the random scan plots, we must first ensure that the mass of $h_{2}$ can conform to the experimental constraints, because the mass determination for SM-like Higgs is already very accurate. In Fig.7 (a), we chose 1 GeV theory uncertainty for the SM-like Higgs. For $h_{1}$, we can choose the theory uncertainty up to 3 GeV, because CMS only shows that there is a 2.8 $\sigma$ excess at 95.3 GeV CMS1 ; CMS5 ; CMS9 , which is not exactly observed. In Fig 7 (b), the $(\mu_{NP}^{\gamma\gamma},\mu_{NP}^{bb})$ plane, we can clearly see that with the parameter space of Tab.2, most points can explain the di-photon excess and the $b\bar{b}$ excess.
@@ V Conclusion
In this paper we introduced CPV in the $\mu\nu$SSM, which lead CP-even Higgs sector mixed with CP-odd Higgs sector. We also analyzed a excess in the di-photon decay mode at $\sim$ 95 GeV as reported by ATLAS and CMS, together with a $\sim$ 2 $\sigma$ excess at LEP in the same mass range. The mixing and CPV are used to produce the lightest Higgs boson mass around 95 GeV, and the next-to-lightest Higgs boson mass around 125 GeV is the so-called SM-like Higgs. The lightest Higgs boson can explain an excess of $\gamma\gamma$ events at $\sim$ 96 GeV as reported by CMS.
In the numerical part, we find a suitable parameter spaces, based on which we show the properties of the lightest and next-to-lightest Higgs boson. We found it very easy get both $\mu_{NP}^{bb}$ and $\mu_{NP}^{\gamma\gamma}$ to reach the central values of their respective experiments at the same time. We also analyze the influence of relevant parameters and CPV phases on the signal strengths or Higgs masses, and improve the signal strengths as much as possible while ensuring that the SM-like Higgs meets the experimental constraints and the “di-photon excess” is around 95 GeV.
@@ Acknowledgements.
The work has been supported by the National Natural Science Foundation of China (NNSFC) with Grants No. 12075074, No. 12235008, No. 11705045, No. 11535002, Hebei Natural Science Foundation with Grant No. A2022201017, No. A2023201041, Natural Science Foundation of Guangxi Autonomous Region with Grant No. 2022GXNSFDA035068, the youth top-notch talent support program of the Hebei Province.
Appendix A Msss matyices

### 06 Mar 2024 v2

@@ III Excess at 95 GeV
The value for $\mu_{LEP}^{bb}$ can be found in 98LEP1 ; 98LHC with the method introduced in 98LEP2 . $h_{1}$ is the Higgs which has mass range around $\sim$ 96 GeV, and $h_{2}$ is the 125 GeV Higgs boson in the following. In the framework of the $\mu$-from-$\nu$ SSM we use $\mu_{NP}^{bb}$ to describe the signal strength; the expression for $\mu_{NP}^{bb}$ can be approximated as 98LEP3
@@
One can find the SM branching ratios $Br^{SM}$ in Ref. PDG1 , and $\Gamma$ is the decay widths. The couplings are normalized to the SM prediction of a Higgs boson of the same mass. $C_{h_{1}}$ is the coupling of $h_{1}$ and gauge boson, and $C_{h_{1}u\bar{u}}$ and $C_{h_{1}d\bar{d}}$ are the couplings of $h_{1}$ and up- and down-type quarks. The normalized couplings are given as
$C_{h_{i}d\bar{d}}={{Z_{H}^{i1}}\over{\cos\beta}},\qquad C_{h_{i}u\bar{u}}={{Z_{H}^{i2}}\over{\sin\beta}},\qquad C_{h_{i}VV}=Z_{H}^{i1}\cos\beta+Z_{H}^{i2}\sin\beta.$ (30)
@@
$\mu_{\gamma\gamma}^{CMS}={{\sigma(gg\rightarrow h_{1}\rightarrow\gamma\gamma)}\over{\sigma^{SM}(gg\rightarrow h_{1}\rightarrow\gamma\gamma)}}=0.6\pm 0.2.$ (31)
Recently, ATLAS reported their new results at 95.4 GeV based on the full Run 2 dataset 95ATLAS , the “diphoton excess” with a signal strength of
@@
Meanwhile, the corresponding CMS result for the “diphoton excess” is given by CMS4
@@
In this work, the approximation of the diphoton rate of the $h_{1}$ can written as 98LEP1 ; 98LEP3
@@
$\approx{{|C_{h_{1}u\bar{u}}|^{2}\times|C_{h_{1}\gamma\gamma}|^{2}}\over{|C_{h_{1}d\bar{d}}|^{2}(Br_{h_{1}\rightarrow b\bar{b}}^{\text{SM}}+Br_{h_{1}\rightarrow\tau\bar{\tau}}^{\text{SM}})+|C_{h_{1}u\bar{u}}|^{2}(Br_{h_{1}\rightarrow gg}^{\text{SM}}+Br_{h_{1}\rightarrow c\bar{c}}^{\text{SM}})}}.$
@@
$|C_{h_{1}\gamma\gamma}|^{2}={{|{4\over 3}C_{h_{1}t\bar{t}}A_{1/2}(\tau_{t})+C_{h_{1}VV}A_{1}(\tau_{W})|^{2}}\over{|{4\over 3}A_{1/2}(\tau_{t})+A_{1}(\tau_{W})|^{2}}}$ (36)
with $\tau_{t}={{m_{h_{1}}^{2}}\over{4m_{t}^{2}}}<1$ and $\tau_{t}={{m_{h_{1}}^{2}}\over{4m_{W}^{2}}}<1$. The form factors $A_{1/2}$ and $A_{1}$ are given by Chrr
@@
By using Eqs. (29) and (35), we calculate the two signal strengths.
@@ IV Numerical Results
$\lambda,\quad\tan\beta,\quad\kappa,\quad A_{t},\quad A_{\lambda},\quad A_{\kappa},\quad A_{b},\quad\upsilon_{\nu^{c}}.$ (39)
We take $\upsilon_{\nu_{1}^{c}}=\upsilon_{\nu_{2}^{c}}=\upsilon_{\nu_{3}^{c}}$, and we have defined
@@ IV.1 Mass and coupling
Figure 1: Values for $m_{h_{2}}$ versus $m_{h_{2}}$ in (a) and the normalized couplings $|C_{h_{2}VV}|$ versus $|C_{h_{1}VV}|$ in (b). The input parameters are in Table 1.
In Fig. 1 (a), we can see that, when $h_{1}$ is near 96 GeV, $h_{2}$ is closer to 125 GeV with the increase of $\tan\beta$. Although both $h_{1}$ and $h_{2}$ can conform to the experimental mass range in our parameter space, if we assume a theory uncertainty of up to 3 GeV 98LEP3 , the parameter range will be larger. One can clear see that $|C_{h_{2}VV}|$ is much larger than $|C_{h_{1}VV}|$ in Fig. 1 (b); the LHC measurements of the SM-like Higgs bosons couplings to fermions and massive gauge bosons are still not very precise98LEP3 ; LHC-couplings . If in the future some collider can measure these couplings to the percent level, then we can choose a more reasonable parameter space.
Figure 2: Values for $m_{h_{1}}$ versus the parameters $\upsilon_{\nu^{c}}$ in (a) and $\kappa$ in (c) and values for $m_{h_{2}}$ versus the parameters $\upsilon_{\nu^{c}}$ in (b) and $\kappa$ in (d). We take $\tan\beta=5$ and other parameters in Table 1.
In Fig.2, left column, we can see that $m_{h_{1}}$ is very sensitive to $\upsilon_{\nu^{c}}$. As $\upsilon_{\nu^{c}}$ becomes larger, $m_{h_{1}}$ will rapidly become smaller; on the contrary, $m_{h_{2}}$ will slowly become larger. In order for $m_{h_{1}}$ to be around 95.4 GeV, the value of $\upsilon_{\nu^{c}}$ will not have a wide range. Similar to $\upsilon_{\nu^{c}}$, $m_{h_{1}}$ is also very sensitive to $\kappa$, and with the increase of $\lambda$, a larger $\kappa$ value can be taken to keep $m_{h_{1}}$ near 95.4 GeV. On the contrary, $\kappa$ has little effect on $m_{h_{2}}$; especially when $\lambda$=0.085 or 0.087, $m_{h_{2}}$ will be very stable.
@@
Figure 3: The component of $h_{1}$ (a) and $h_{2}$ (b). The input parameters are in Table 1, and we take $\tan\beta=4.31$.
In Fig. 3, we have showed the component of $h_{1}$ and $h_{2}$; for $h_{1}$, $CP$-odd $H_{u}$ component and $CP$-even right-handed sneutrinos component are the main components, and with the increase of $\lambda$, $CP$-odd $H_{d}$ component will become smaller, while the $CP$-even right-handed sneutrino component will become larger. While for $h_{2}$, $CP$-even $H_{u}$ is the main component, as $\lambda$ increases, the $CP$-even $H_{u}$ component will increase, the $CP$-even $H_{d}$ component will gradually decrease, and the $CP$-odd $H_{d}$ component will gradually increase.
Figure 4: The correlation of $h_{1}$ and the CPV phases in (a) (c) (e) (g); the gray area represents the experimental error of 3 GeV. The correlation of $h_{2}$ and the CPV phases in (b) (d) (f) (h); the gray area is the theory uncertainty of 1 GeV.
Then, in Fig. 4, we analyze the correlation between the Higgs masses and the CPV phases. For the first row, we take $\lambda=0.09,\tan\beta=6,$ and $\kappa=0.315,$ for the second row, $\lambda=0.095,\tan\beta=4.5,$ and $\kappa=0.315,$ for the third row, $\lambda=0.098,\tan\beta=4.3,$ and $\kappa=0.32$, for the last row, $\lambda=0.085,\tan\beta=3,$ and $\kappa=0.04$. Other parameters remain the same as Table 1. We should remark that there are many CPV phase values that can be constrained by the Higgs masses in Fig.7 (a), and here we select only one of them. The signal strengths and the Higgs masses do not increase or decrease all the time, but oscillate like a sine function. In Fig. 4, the first row, when $\phi_{\lambda}$ grows from 4 to 5, the mass range of the “diphoton excess” can fit the experimental constraints, and the SM-like Higgs mass can be kept around 125 GeV.
@@ IV.2 Signal strengths
Figure 5: The left column shows that the signal strength $\mu_{NP}^{bb}$ varies with $\upsilon_{\nu^{c}}$ (a) and $\kappa$ (c); the right column shows that the signal strength $\mu_{NP}^{\gamma\gamma}$ varies with $\upsilon_{\nu^{c}}$ (b) and $\kappa$ (d). We take $\lambda=0.09$ and other parameters as in Table 1.
@@
The effect of CPV phases is also very obvious. Here, we take $\phi_{A_{t}}$ as an example, we take $\phi_{A_{t}}$ from $-\pi$ to $\pi$. In the first row in Fig. 6, we can see that the minimum value of $h_{1}$ appears to be around 96.2 GeV and varies periodically with $\phi_{A_{t}}$. The maximum value of $h_{1}$ does not exceed 1 GeV larger than the minimum value. Meanwhile, the highest point of $h_{2}$ is at 125.3 GeV, with a maximum value 1.5GeV larger than the minimum value.
The impact of $\phi_{A_{t}}$ on the normalized couplings is relatively small. In the second row in Fig. 6, we can see that the peak of $|C_{h_{1}VV}|$ is close to 0.313, and the highest point of $|C_{h_{2}VV}|$ is slightly less than 0.856. This implies that, if the ”95.4 excess” is a new particle, then its coupling with gauge bosons should be much smaller than the coupling of the SM-like Higgs with gauge bosons.
Figure 6: The Higgs masses vary with the phase $\phi_{A_{t}}$ in (a)-(b); the correlation of couplings and the phase $\phi_{A_{t}}$ in (c)-(d); and (e)-(f) are the correlation of signal strengths and the phase $\phi_{A_{t}}$. Here, we take $\tan\beta$=4.5, $\lambda$=0.086, and $\phi_{A_{t}}=[-\pi,\pi]$, and other parameters are the same as in Table 1.
@@
Let us remark that, in the random scan plots, we must first ensure that the mass of $h_{2}$ can conform to the experimental constraints, because the mass determination for SM-like Higgs is already very accurate. In Fig.7 (a), we chose 1 GeV theory uncertainty for the SM-like Higgs. For $h_{1}$, we can choose the theory uncertainty up to 3 GeV, because CMS shows only that there is a 2.8 $\sigma$ excess at 95.3 GeV CMS1 ; CMS5 ; CMS9 , which is not exactly observed. In Fig 7. (b), the $(\mu_{NP}^{\gamma\gamma},\mu_{NP}^{bb})$ plane, we can clearly see that with the parameter space of Table 2, most points can explain the dphoton excess and the $b\bar{b}$ excess.
@@ V Conclusion
In this paper we introduced CPV in the $\mu$-from-$\nu$SSM, which leads to $CP$-even Higgs sector mixed with $CP$-odd Higgs sector. We also analyzed an excess in the diphoton decay mode at $\sim$ 95 GeV as reported by ATLAS and CMS, together with a $\sim$ 2 $\sigma$ excess at LEP in the same mass range. The mixing and CPV are used to produce the lightest Higgs boson mass around 95 GeV, and the next-to-lightest Higgs boson mass around 125 GeV is the so-called SM-like Higgs. The lightest Higgs boson can explain an excess of $\gamma\gamma$ events at $\sim$ 96 GeV as reported by CMS.
In the numerical part, we find a suitable parameter space, based on which we show the properties of the lightest and next-to-lightest Higgs boson. We found it very easy get both $\mu_{NP}^{bb}$ and $\mu_{NP}^{\gamma\gamma}$ to reach the central values of their respective experiments at the same time. We also analyze the influence of relevant parameters and CPV phases on the signal strengths or Higgs masses, and improve the signal strengths as much as possible while ensuring that the SM-like Higgs meets the experimental constraints and the “diphoton excess” is around 95 GeV.
@@ Acknowledgements.
The work has been supported by the National Natural Science Foundation of China (NNSFC) with Grants No. 12075074, No. 12235008, No. 11705045, and No. 11535002, Hebei Natural Science Foundation with Grants No. A2022201017 and No. A2023201041, Natural Science Foundation of Guangxi Autonomous Region with Grant No. 2022GXNSFDA035068, and the youth top-notch talent support program of the Hebei Province.
Appendix A mass matrices